OpenStudy (anonymous):

100 plus the product of a number and -2 equals 50. What is the number? (Be sure to type only the number in the answer blank. No variables or words.)

OpenStudy (defiance):

Okay, I like to reverse the problem. 100 + ? - 2 = 50 We see the number went down, so the answer is a negative.

OpenStudy (defiance):

So, 50 + 2 = ?

OpenStudy (defiance):

100 - 52 = ? ? = 48

OpenStudy (anonymous):

52

OpenStudy (defiance):

Mhm

OpenStudy (defiance):

Answer is 48

OpenStudy (defiance):

At the end.

OpenStudy (defiance):

wait

OpenStudy (defiance):

Accutally, you're right, lol.

OpenStudy (defiance):

Sorry, im getting mixed,.

OpenStudy (defiance):

It's -52

OpenStudy (anonymous):

Wait im confused

OpenStudy (defiance):

Okay.

OpenStudy (defiance):

100 - ? - 2 = 50. This ? = 48

OpenStudy (defiance):

The answr is -48

OpenStudy (anonymous):

Ok u understand thanj you

OpenStudy (texaschic101):

100 + (-2n) = 50 100 - 2n = 50 -2n = 50 - 100 -2n = -50 n = -50/-2 n = 25

OpenStudy (texaschic101):

the product is a result of multiplication.....so the product of -2 and a number is -2n. So we have 100 + (-2n) = 50